carol needs to travel from her home in Collinsville to stoney point. She can travel 25 miles via a rural route at 50mph or a longer distance at 70mph along the interstates. The eastern portions of the interstate is 5 miles less than the northern portion of the interstate.
how long(time) will it take carol to drive the rural route?
how long (miles) will it take Carol to drive the interstate route?
How long (time) will it take Carol to drive the interstate route?

Found 2 solutions by addingup, ikleyn:
Answer by addingup(3677)

50 miles = 1 hour
25 miles = 50/2 = 1 hour/2 = 1/2 hour
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
"The eastern portions of the interstate is 5 miles less than the northern portion of the interstate"
How long is the interstate? Or how long is the northern portion?

Answer by ikleyn(52803)

.

As I understand the condition, it represents (in encrypted form) the following information:

    There are two legs of the right angled triangle - interstate highways,

    the eastern portion, and the northern portion. 

    One leg is 5 miles shorter than the other.

    The hypotenuse is 25 miles - the rural route.

I don't like very much this way of presenting problems, when the solution is to de-encrypt the authors words.

So, I will not solve it farther. My work on de-encrypting is just enough.

I only want to add that from my de-encrypting it follows immediately that the triangle is (3-4-5) right-angled

triangle with the legs of 15 and 20 miles and the hypotenuse of 25 miles.


From this point, you can easily complete the solution on your own.